when solving a linear system such as this one: 9x-7y=31
-9x+3y=-39
how do you do it? I know its a simple problem but i'm confused on what to subtract or add.

Answer by Alan3354(69443)

when solving a linear system such as this one:
9x-7y=31
-9x+3y=-39
how do you do it?
----------
Since both x terms have a coefficient of 9,
9x-7y=31
-9x+3y=-39
--------------- Add
-4y = -8
etc